OpenStudy (anonymous):

What is the sum of a 38–term arithmetic sequence where the first term is 14 and the last term is 154? a) 2660 b)2850 c)2926 d)3192

OpenStudy (anonymous):

Well first you have to find out what the difference between terms is. Do you know how to do that?

OpenStudy (anonymous):

O wait no you don't sorry didnt read your question carefully.

OpenStudy (anonymous):

You don't have to.

OpenStudy (anonymous):

do you use the arithmetic sequence formula=a1+(n-1)d substitute the values and find d??

OpenStudy (anonymous):

Well, I'd rather have you do it the following way. Let's focus on how to get the sum. We have an arithmetic sequence, with a term \[t_n\] being the nth term. Thus the thing you're looking for is: \[t_1+t_2+...t_34\]

OpenStudy (anonymous):

Correction: \[t_{34}\]

OpenStudy (anonymous):

Good so far?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Now I propose we do this: Pair the first and last terms, then second and second to last, etc, etc. This is the commutative and associative property of addition, so it's got to be valid: \[t_1+t_{34}+t_2+t_{33}...=\] \[(t_1+t_{34})+(t_2+t_{33})...\]

OpenStudy (anonymous):

Good?

OpenStudy (anonymous):

Um..yes

OpenStudy (anonymous):

Now let's say you have the mean m such that \[m = \frac{t_1+t_{34}}{2}\]

OpenStudy (anonymous):

Okay. Now let's say I look at \[\frac{t_2+t_{33}}{2}\]

OpenStudy (anonymous):

Use the formula you mentioned earlier with difference d between terms to substitute t2 and t33 with t1 and t34

OpenStudy (anonymous):

Well?

OpenStudy (anonymous):

so the its 154+14/2?

OpenStudy (anonymous):

No, no. Let's go symbolic. Replace t2 and t33 using only t1, t34, and d.

OpenStudy (anonymous):

huh?ok sorry but now im really lost..

OpenStudy (anonymous):

Okay. Do you agree that \[t_2 = t_1-d\]and\[t_33=t34-d\], right?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

wait no made a couple typos: \[t_2 = t_1+d\]and\[t_{33}=t_{34}-d\]

OpenStudy (anonymous):

There. Now, going back to our little example of \[\frac{t_2+t_{33}}{2}\]

OpenStudy (anonymous):

Replace t2 and t33...

OpenStudy (anonymous):

Well, what do you get?

OpenStudy (anonymous):

2660?

OpenStudy (anonymous):

Look, I'm trying to lead you to the answer. Stay symbolic. K, I'll just show you. So we had: \[m=\frac{t_1+t_{34}}{2}\]What I wanted you to do was: \[\frac{t_2+t_{33}}{2}=\frac{t_1+d+t_{34}-d}{2} = \frac{t_1+t_{34}}{2}\]agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then, you should see that \[\frac{t_3+t_{32}}{2}\] also equals m

OpenStudy (anonymous):

By using \[t_3 = t_2 + d\]and\[t_{32}=t_{33}-d\]and then you get\[\frac{t_{33}+d-d+t_2}{2}=\frac{t_{33}+t_2}{2} =\frac{t_{34}+t_1}{2} = m\]

OpenStudy (anonymous):

Agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. Then, going back to my sequence... \[(t_1+t_{34})+(t_2+t_{33})...\]

OpenStudy (anonymous):

Which is still the original sum you're looking for, I do the following, multiplying by 1. \[(t_1+t_{34})+(t_2+t_{33})...=\]\[1*((t_1+t_{34})+(t_2+t_{33})...)=\]\[(2/2)*((t_1+t_{34})+(t_2+t_{33})...)=\]\[2*((t_1+t_{34})/2+(t_2+t_{33})/2...)=\]\[2*(m+m...)=\]

OpenStudy (anonymous):

Do you know how many m's there will be summed together?

OpenStudy (anonymous):

look the question has 4 possible answers a)2660 b)2850 c)2926 and d)3192

OpenStudy (anonymous):

Well, I could simply answer for you but this proof is much more important; that is, only if you want to learn...

OpenStudy (anonymous):

Okay. Anyway. Notice that there were 34 numbers. I PAIRED them. How many PAIRS are there - each pair turned into an "m"

OpenStudy (anonymous):

17?

OpenStudy (anonymous):

Yes! 17! Bettern known as 34/2! \[2*(m+m+m...)=2*(m\frac{34}{2})=34 m\]

OpenStudy (anonymous):

Remember, this was originally \[t_1+t_2...+t_{34}\]

OpenStudy (anonymous):

So there's your answer. Take the mean of t1 and t34, and then multiply by the number of terms, 34!

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

Well?

OpenStudy (anonymous):

but how do u get the mean of t1 and t34?

OpenStudy (anonymous):

I'm really confused -.-

OpenStudy (anonymous):

t1 is the first term, and t34 is the last term. How's it that hard?

OpenStudy (anonymous):

In fact, for any arithmetic sequence of length n, the sum is always \[n*mean = n\frac{t_1+t_n}{2}\]

OpenStudy (anonymous):

I still dont get it!!! plzz help!

OpenStudy (anonymous):

Okay. Let's try a different approach... Imagine you have a 4-term arithmetic series: \[t_1+t_2+t_3+t_4\]

OpenStudy (anonymous):

You know that between each term there's a difference d.

OpenStudy (anonymous):

Forget it..I got it thnxz anyways

OpenStudy (anonymous):

The answer is 3192

OpenStudy (anonymous):

You use the formula Sn=n/2(a1+an)

OpenStudy (anonymous):

Thats what I proved!!!

OpenStudy (anonymous):

Where n is 38, a1 is 14 and an is 154

OpenStudy (anonymous):

What I said earlier, verbatim: In fact, for any arithmetic sequence of length n, the sum is always \[n∗mean=n\frac{t_1+t_2}{2}\]

OpenStudy (anonymous):

t_n not t_2**

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

Your equation for Sn is the exact same, but uses a instead of t... But fine. If you insist on memorizing a formula, then you're ok, right?

OpenStudy (anonymous):

Yah I'm ok i hav a pictographic memory I can't help but memorize

OpenStudy (anonymous):

Pretty sure that's not proven to exist but w/e.

OpenStudy (anonymous):

Whatever...

OpenStudy (anonymous):

K anyway remember to close the question!